In lecture we saw the Cournot competition model for two firms with the same cost function. Now, we are going to consider asymmetric cost functions. Assume that demand for a good is given by
1
Expert's answer
2020-10-28T09:15:01-0400

Reaction function of firm1:

"2Q{_1}+Q{_2}=\frac{a-C{_2}}{b}"       


Reaction function of firm2:

 "Q{_1} - 2Q{_2} = \frac{(a - C2)}{b}"

Explanation:

Demand function:

P = a - bQ

Q = Q1 + Q2

Thus,

P = a - b(Q1 + Q2)

MC1 = C1

TC = C1Q1

MC2 = C2

TC2 = C2Q2


Profit of the firm 1:

Profit1 = TR1 - TC1

= PQ1 - C1Q1

= [a - b(Q1 + Q2)]Q1 - C1Q1

=aQ1 - b(Q1)2 - bQ1Q2 - C1Q1


dProfit1/dQ1 = a - 2bQ1 - bQ2 - C1

dProfit1/dQ1 = 0

a - 2bQ1 - bQ2 - C1 = 0

2Q1 + Q2 = (a - C1)/b              ... (Reaction function of firm 1)


This reaction function shows that optimal quantity of the firm 1 depends on the optimal quantity of the firm two.

Putting the given information in the reaction function of firm 1 as follows:


"2Q{_1} + Q{_2} = \frac{(a - C1)}{b }"   

          

"2Q{_1} + 100 = \frac{(4 - 2)}{0.01}"

2Q1 = 200

Q1 = 100


Thus, if firm 2 produces the 100 units then the firm 1 will produce 100 units.


Profit of the firm 2:

Profit2 = TR2 - TC2

= PQ2 - C2Q2

= [a - b(Q1 + Q2)]Q2 - C2Q2

=aQ2 - bQ1Q2 - b(Q2)2 - C2Q2


dProfit2/dQ2 = a - bQ1 - 2bQ2 - C2

dProfit2/dQ2 = 0

a - bQ1 - 2bQ2 - C2 = 0

Q1 - 2Q2 = (a - C2)/b           ... (Reaction function of firm 2)